If consumption function of an economy is C=40+0.9Y.What will be the investment multiplier.


1
Expert's answer
2020-11-30T16:30:22-0500

"\bold {Answer}"

Investment Multiplier, "K = 10"


"\bold {Solution}"


Invesent Multiplier, "k = \dfrac {∆Y}{∆I}"


It is the ratio of the change in national income induced by a $1 change in investment.

Given the consumption function,

"C = 40 + 0.9Y"

0.9 is the marginal propensity to consume (MPC)

Since, MPC + MPS = 1

=> MPS = 1 - MPC

= 1 - 0.9

= 0.1


Now, "K = \dfrac {∆Y}{∆I} = \dfrac {1}{MPS}"


"= \dfrac {1}{1-0.9}"


"=\dfrac {1}{0.1}"

"= \bold {10}"


Therefore, the investment multiplier is 10.